OpenStudy (anonymous):

y=1/2 cos (3x+3.14/2) need help not sure how to do it.

OpenStudy (mary.rojas):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Data_LG2 you can't distribute like that

OpenStudy (anonymous):

why?

jimthompson5910 (jim_thompson5910):

because cos isn't being multiplied out front cos is a function

OpenStudy (anonymous):

oh ok, sorry.. i'll try again

OpenStudy (anonymous):

@jim_thompson5910 do we need to use identities here?

OpenStudy (anonymous):

yes but that is not what i'm suppose to do. I forgot to look at the directions for this problem I just wrote it down. I think it is a vertical shift. and I have to graph it also. sorry I should have told y'all.

jimthompson5910 (jim_thompson5910):

you could use identities, but acegirl may just want to graph this or figure out the period/phase shift

OpenStudy (anonymous):

oh transformations????, i thought we need to solve it

jimthompson5910 (jim_thompson5910):

In general y = A*cos(Bx - C) + D has an amplitude of |A| has a period of T = 2pi/B has a phase shift of C/B (this is the horizontal shift) and D takes care of the vertical shift

OpenStudy (anonymous):

@jim_thompson5910 that's what i'm working with

jimthompson5910 (jim_thompson5910):

so in this case, what are A, B, C, D?

OpenStudy (anonymous):

amp=1/2 period=2pi/B=2pi/3 p.s.=C/B=pi/2/3 not sure how to work that out

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

amplitude and period are correct

jimthompson5910 (jim_thompson5910):

phase shift C/B = (-pi/2)/3 = (-pi/2)*(1/3) = -pi/6 so this means you shift the graph pi/6 units to the left

jimthompson5910 (jim_thompson5910):

the original graph of cos(x) has not been shifted vertically to get 1/2 cos (3x+pi/2) since D = 0

OpenStudy (anonymous):

where did you get the negative pi from?

jimthompson5910 (jim_thompson5910):

well notice the form I'm working with is y = A*cos(Bx - C) + D pay particular attention to the -C

OpenStudy (anonymous):

okay

jimthompson5910 (jim_thompson5910):

y = 1/2 cos (3x+pi/2) is the same as y = 1/2 cos ( 3x- (-pi/2) ) + 0 now compare that to y = A*cos(Bx - C) + D and we see that A = 1/2 B = 3 C = -pi/2 D = 0

OpenStudy (anonymous):

okay

jimthompson5910 (jim_thompson5910):

hopefully all that makes sense

OpenStudy (anonymous):

so is the phase shift pi/6?

jimthompson5910 (jim_thompson5910):

units to the left

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

If C is negative, you shift left If C is positive, you shift right

OpenStudy (anonymous):

yes it does and thank you. @jim_thompson5910

jimthompson5910 (jim_thompson5910):

ok glad it does, yw